solve by substitution
5x + 3y = 8
-7x + y = 20

Answer by mananth(16946)

5 x + 3 y = 8 .............1
1 y = 20 + 7 x ---------2
/ 1 =
y = 20 + 7 x
Substitute y in (1)
5 x + 3 ( 20 + 7 x ) = 8
5 x + 60 + 21 x = 8
5 x + 21 x = 8 -60
26 x = -52
/ 26
x= -2
Plug the value of x in (1)
5 x + 3 y = 8
5 * -2 + 3 y = 8
-10 + 3 y = 8
3 y = 18
/ 3
y= 6
m.ananth@hotmail.ca
CHECK
1 y = 20 + 7 x
6 = 20 + -14